OpenStudy (anonymous):

The straight line y=2x+k meets the parabola y=x^(2)+6x-5 at two distinct points A and B. (a) Find the range of values of k. (b) Suppose the coordinates of A are (2,11). (i) Find the value of k. (ii) Find the coordinates of B.

hartnn (hartnn):

can you solve y=2x+5 and y=x^(2)+6x-5 simultaneously ?

hartnn (hartnn):

put y=2x+5 in y=x^(2)+6x-5

OpenStudy (anonymous):

that should be 2x+k not 2x+5

hartnn (hartnn):

yes.

hartnn (hartnn):

so , what u get ?

OpenStudy (anonymous):

no, that should be 2x+k............. i got the answer wrongly, the answer is k>-9 but my answer is k>-16...

OpenStudy (dmezzullo):

can you show your work?

hartnn (hartnn):

yes, that was typo...i meant 2x+k only.

OpenStudy (anonymous):

ohh, wait.....maybe i get the wrong question..................

hartnn (hartnn):

so, you already knew that you need to make discriminat >0 for the quadratic you gonna get ?

hartnn (hartnn):

*discriminant damn

OpenStudy (anonymous):

yup, i got the answer of part a now:)

hartnn (hartnn):

so, A = (2,11) which makes x=2, can u get 'k' now ?

OpenStudy (anonymous):

but part b..... sub x=2 and y=11 into the equation y=2x+k. 11=2(2)+k k=11/4

hartnn (hartnn):

11 = 4+k k= 11-4 =...

OpenStudy (anonymous):

nani 0.0?! oh my god

hartnn (hartnn):

happens with all of us :P

hartnn (hartnn):

what about B ?

hartnn (hartnn):

know how to find it ?

OpenStudy (anonymous):

umm.....

OpenStudy (anonymous):

i got many decimal places......

hartnn (hartnn):

x^2+6x-5 = 2x+7 solve this quadratic.

hartnn (hartnn):

one root = 2, other root =... ?

OpenStudy (anonymous):

x=2 or x=-6 y=11 or y=-5 ??

hartnn (hartnn):

\(\huge \color {red}\checkmark\)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^